OpenStudy (precal):

Related Rates A spherical balloon is expanding at a rate of 60 pi cubic inches per second. How fast is the surface area of the balloon expanding when the radius of the balloon is 4 inches?

OpenStudy (precal):

Am I solving for dr/dt?

zepdrix (zepdrix):

Solving for dA/dt it looks like. Where A is the surface area of a sphere.

zepdrix (zepdrix):

Volume of a sphere, \[\Large\rm V=\frac{4}{3}\pi r^3\]Surface Area of a sphere,\[\Large\rm A=4\pi r^2\]

OpenStudy (precal):

ok I was not using both of those

OpenStudy (precal):

the cubic inches per second threw me off

zepdrix (zepdrix):

So we're given \(\Large\rm r=4\), and \(\Large\rm dV/dt=60\), yes? Or does expanding mean the rate at which it's moving outward? radius change? Oh no no, the units are cubic, so they're talking about volume change, ok ok ok.

OpenStudy (precal):

yes 60 pi for dv/dt

OpenStudy (precal):

\[60 \pi\]

OpenStudy (precal):

I am confused on this one

OpenStudy (precal):

I was using surface area of a sphere

zepdrix (zepdrix):

So ummm let's see if we can figure this out. Differentiating our Volume function, with respect to time,\[\Large\rm \frac{dV}{dt}=4\pi r^2\frac{dr}{dt}\]Does that look ok?

OpenStudy (precal):

yes that part looks ok

zepdrix (zepdrix):

So in order to find dA/dt, it looks like we'll need to deal with this dr/dt first. So our first step, yes, is to solve for dr/dt using the information they gave us.

OpenStudy (precal):

ok let me do that

OpenStudy (precal):

15/16

OpenStudy (precal):

or .9375

zepdrix (zepdrix):

Ok great.\[\Large\rm \frac{dr}{dt}=\frac{15}{16}\]

zepdrix (zepdrix):

Let's jump over to our surface area function now, we're done using the volume formula, it gave us what we needed.

OpenStudy (precal):

30 pi

OpenStudy (precal):

inches^2 per second

zepdrix (zepdrix):

Ooo yay good job \c:/

OpenStudy (precal):

Thank you